### v2

@@ 4.6 User study
Results of Test 2 are reported on the right of Fig. 5. It is worth noting that these ratings are closely linked to both the scan quality and the fidelity of the mouth representation. Despite the scan quality not being optimal, as depicted in Fig. 5, our results indicate that we can achieve a good level of realism in the animated scans. The interface for Test 1 and Test 2 can be accessed in the supplementary material Sec. 6.7.
@@ 5 Conclusions and Limitations
While ScanTalk exhibits commendable performance in lip motion synthesis, we refrained from incorporating expressions due to the persistently limited availability of such data. Exploring how a similar framework to ScanTalk would handle these additional modalities could be an engaging avenue for future research. Although ScanTalk can animate unregistered meshes, its current training process necessitates meshes with a common topology within each sequence. To overcome this limitation, a fully unsupervised training strategy with a loss function that does not rely on point-wise correspondence, yet captures small geometric displacements, could be a promising direction for further exploration.
@@ 6.4 Datasets
| VOCAset | BIWI | Multiface | |
@@ 6.5 Mesh encoding
Several encoding strategies for geometry are feasible; however, with our experimentation we saw that encoding vertex positions provides the optimal and most intuitive approach. When omitting mesh encoding and directly feeding the BiLSTM output to the decoder, the mesh signal remains constant across frames, leading to a static facial expression as the decoder lacks spatial awareness of the mouth’s location. Incorporating normals alongside positions fails to enhance results, as precomputed operators already furnish adequate orientation information. Additionally, adopting the Heat Kernel Signature (HKS), as suggested in the DiffusionNet framework, does not yield improvements in results. In Fig. 8, we present the per-vertex norm of features derived from the DiffusioNet Encoder for both training and testing meshes.

### v3

@@ 4.6 User study
Results of Test 2 are reported on the right of Fig. 6. These ratings are closely linked to both the scan quality and the fidelity of the mouth representation. Despite the scan quality not being optimal, as depicted in Fig. 6, our results indicate that we can achieve a good level of realism in the animated scans.
@@ 5 Conclusions
Acknowledgments
This work is supported by the ANR project Human4D ANR-19-CE23-0020 and by the IRP CNRS project GeoGen3DHuman. This work was also partially supported by “Partenariato FAIR (Future Artificial Intelligence Research) - PE00000013, CUP J33C22002830006" funded by NextGenerationEU through the italian MUR within NRRP, project DL-MIG. This work was also partially funded by the ministerial decree n.352 of the 9th April 2022 NextGenerationEU through the italian MUR within NRRP. This work was also partially supported by Fédération de Recherche Mathématique des Hauts-de-France (FMHF, FR2037 du CNRS).
@@ 6.4 Datasets
VOCAset | BIWI6 | Multiface | |
@@ 6.5 Mesh encoding
Several encoding strategies for geometry are feasible; however, with our experimentation we saw that encoding vertex positions provides the optimal and most intuitive approach. When omitting mesh encoding and directly feeding the BiLSTM output to the decoder, the mesh signal remains constant across frames, leading to a static facial expression as the decoder lacks spatial awareness of the mouth’s location. Incorporating normals alongside positions fails to enhance results, as precomputed operators already furnish adequate orientation information. Additionally, adopting the Heat Kernel Signature (HKS), as suggested in the DiffusionNet framework, does not yield improvements in results. In Fig. 9, we present the per-vertex norm of features derived from the DiffusionNet Encoder for both training and testing meshes.